OpenStudy (anonymous):

sqrt11/(50m) how do you do this radical to see if its in simplified form? Need help please

OpenStudy (kkutie7):

\[\sqrt{\frac{ 11 }{ 50 }}\] right? for a start you can split it up like this \[\frac{\sqrt{11}}{\sqrt{50}}\]

OpenStudy (anonymous):

ok yea

OpenStudy (anonymous):

then what do I do

OpenStudy (kkutie7):

well you can't break down 11 but you might be able to break down 50 let me show you.

OpenStudy (anonymous):

no I have not im trying to learn all this stuff gets confusing to me

OpenStudy (kkutie7):

oh I need to edit that I messed up let me fix it so i can explain =)

OpenStudy (kkutie7):

 Ok so what I did here was break down 50 into its multiples 25*2=50 you can break down 2 any more but you can 25 5*5= 25 so that is how I got that

OpenStudy (kkutie7):

 we know that the sqrt of 25 is 5 so you can pull it out writing the fraction like this \[\frac{\sqrt{11}}{5\sqrt{2}}\]

OpenStudy (kkutie7):

you can rewrite it any way you want here is another way \[\frac{\sqrt{\frac{11}{2}}}{5}\]

OpenStudy (kkutie7):

do you understand? If not I'd be happy to answer any of your questions

OpenStudy (anonymous):

a little bit so that means with the 11 its not in simplified form cause there is no way of breaking that down

OpenStudy (ajspeller):

11 is prime. After prime factorization: 11 50 is composite. After prime factorization: 5*5*2 \[\frac{ \sqrt[2]{11} }{ \sqrt[2]{50} }=\frac{ \sqrt[2]{11} }{ \sqrt[2]{2*5*5} }=\frac{ \sqrt[2]{11} }{ 5\sqrt[2]{2} }\] Then you should rationalize ... \[\frac{ \sqrt[2]{11} }{ 5\sqrt[2]{2} }*\frac{ \sqrt[2]{2} }{ \sqrt[2]{2} }=\frac{ \sqrt[2]{11*2} }{ 5\sqrt[2]{2*2}}=\frac{ \sqrt[2]{22} }{ 5\sqrt[2]{4} }=\frac{ \sqrt[2]{22} }{ 5*2 }=\frac{ \sqrt[2]{22} }{ 10 }\]